[Brahms-l] Shift report 20050206 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Feb 06 2005 - 00:01:59 EST
1. Had beam from about 18.45
2. Did runs only for MRS.
MRS@45deg@B2255, tr3:
run 12821: 269k
run 12822: 117k
TPC pedestal run - 12819
Non-TPC pedestal run - 12820
3. Big problem with mainframe L1458B, Bob and Hiro are working on it now

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 6 00:02:18 2005

This archive was generated by hypermail 2.1.8 : Sun Feb 06 2005 - 00:02:40 EST